OpenStudy (anonymous):

If 6c = 4T2 - 50, which of the following could T possibly be?

OpenStudy (anonymous):

T = square root of negative 6 times C plus 50 over 4 T = square root of negative 6 times C minus 50 over 4 T = square root of 6 times C minus 50 over 4 T = square root of 6 times C plus 50 over 4

OpenStudy (anonymous):

the equation tool works well here

OpenStudy (anonymous):

sorry i forgot about that button

OpenStudy (anonymous):

\[t=\sqrt{-6c+50/4}\]

OpenStudy (anonymous):

\[t=\sqrt{-6-50/4}\]

OpenStudy (anonymous):

i meant for the question

OpenStudy (anonymous):

\[t=\sqrt{6c-50/4}\]

OpenStudy (anonymous):

\[t=\sqrt{6c+50/4}\]

OpenStudy (anonymous):

those are A B C and D

OpenStudy (anonymous):

in B there is supposed to be a C next to the 6

OpenStudy (anonymous):

\[6c = 4T^2-50\] \[6c + 50 = 4T^2\] \[(6c+50)/4 = T^2\] \[\sqrt{(6c+50)/4} = T\]

OpenStudy (anonymous):

Sorry it took so long, thought I would just show you each step I took to the answer.and that equation thing is aggravating. :)

OpenStudy (anonymous):

lol its ok thank you so much, you are a big help, this is like the only lesson i dont get in algebra